OpenStudy (anonymous):

A study of one thousand teens found that the number of hours they spend on social networking sites each week is normally distributed with a mean of 10 hours. The population standard deviation is 2 hours. What is the 95% confidence interval for the mean? 9.88−10 hours 10−10.12 hours 9.88−10.12 hours 9.76−10.24 hours

OpenStudy (anonymous):

0.196 is the constant for a 95% confidence interval Lower limit =Mean - (0.196) x\[\frac{ Standard deviation value }{ \sqrt{mean}}\] Upper limit =Mean + (0.196) x \[\frac{ Standard deviation value }{ \sqrt{mean}}\]